Grade 12Mechanics

A motor car covers 1/3rd part of total distance with v1= 10 km/hr, second 1/3rd part with V2= 20km/hr and rest 1/3rd part with V3=60km/he.what is the average speed of the car?( Answers with steps).

Kaviya
8 Years agoGrade 12

1 Answer

Shreya Gautam
ApprovedApproved Tutor Answer8 Years ago
Let total distance covered be S.Time taken = distance covered / speedTime taken in first one third part of the journey= (S/3)/10= S/30Time taken in second one third of the journey= (S/3)/20= S/60Time taken in last one third of the journey= (S/3)/60= S/180Total time = S/30 + S/60 + S/180 = S/18Average speed = total distance/ total time taken= S/(S/18) = 18 km/hr